The energy of electron in excited H-atoms is -3.4eV. Find the angular momentum of electron

Grade:11

1 Answers

Arun
4 years ago
Dear student
 
angular momentum = sqrt (2 me Ke)
 
put the value of m = 9.1 *10^-31 and K = 3.4 * 1.6*10^-19 Jule